# ICon: In-Context Contribution for Automatic Data Selection

### v1

@@ 7 Conclusion
In this paper, we introduced ICon, a novel method that uses ICL to efficiently assess the contribution of individual training samples without human inductive bias. The method is structured into three stages: constructing an assessment set, computing ICon scores, and training a lightweight selection paradigm for scalable data selection. Extensive experiments on three LLMs across 12 benchmarks and 5 pairwise evaluation sets show that models trained on ICon-selected data outperform those trained on the full dataset, achieving superior performance with fewer samples. ICon also significantly outperforms other widely used data selection methods. We further analyzed the optimal data scale, generalization across datasets, and characteristics of high-contribution samples. As data scale increases, performance improves initially but later declines, indicating an optimal data size. ICon generalizes well across datasets, and its high-contribution samples display diversity while maintaining an appropriate level of difficulty for effective learning.
@@
While ICL can be seen as a form of implicit fine-tuning, it captures the impact of individual samples on the model’s behavior, which may not fully replicate the effect of batch training with gradient updates. As a result, some differences in performance may arise. Nevertheless, we believe that ICL provides a useful approximation of sample contribution and can still effectively guide data selection for instruction tuning despite these differences. Furthermore, while our method demonstrates promising results for instruction tuning, we believe it has the potential for broader applicability. For instance, ICon could be adapted to other tasks and domains by modifying the assessment set and the calculation of the ICon score. Exploring these extensions is beyond the scope of this paper and remains a direction for future work.
@@ Appendix A Detailed Evaluation Benchmark for Different Methods
The detailed results for different methods on evaluation benchmarks are shown in Table 5. The methods include Low PPL, Top PPL, Alpagasus, Deita, Superfilter, and ICon. Overall, ICon outperforms existing methods across the evaluated benchmarks.
@@
ICon (Ours) 21.16 29.89 25.52 60.58 30.81 52.30 43.16 46.71 44.80 44.15 55.56 28.42 41.64 67.96 47.09 43.37
Table 5: The detailed results of different methods on the evaluation benchmarks. The datasets AE, TQA, HS, LQA, and WG correspond to AlpacaEval, TruthfulQA, HellaSwag, LogicQA, and WinoGrande, respectively. All models are trained on LLaMA3.1-8B with 7,800 selected samples. The best results are highlighted in bold.
@@ Appendix B Detailed Evaluation Benchmark for Different Data Scales
ICon (1%) 8.31 28.62 18.46 58.47 31.31 55.77 43.11 47.17 49.91 44.51 61.19 23.35 39.37 71.03 48.23 42.91
ICon (5%) 16.21 28.96 22.58 56.91 27.27 53.70 40.51 44.60 47.18 44.23 58.44 24.42 37.75 69.46 46.91 42.09
ICon (10%) 16.08 27.99 22.03 57.86 26.76 53.57 42.63 45.21 45.48 42.81 56.27 26.57 42.01 66.93 46.68 42.08
ICon (15%) 21.16 29.89 25.52 60.58 30.81 52.30 43.16 46.71 44.80 44.15 55.56 28.42 41.64 67.96 47.09 43.37
ICon (20%) 11.95 27.24 19.60 55.07 29.29 53.02 41.71 44.77 43.36 42.08 55.06 26.88 37.36 64.88 44.94 40.66
ICon (25%) 15.63 23.35 19.49 50.84 28.28 51.12 41.26 42.88 42.06 42.13 54.52 28.88 39.77 67.32 45.78 40.43
ICon (30%) 17.72 25.87 21.79 58.57 24.74 52.41 40.66 44.10 41.89 42.32 54.35 26.57 37.63 66.54 44.88 40.77
ICon (50%) 15.00 22.91 18.95 57.75 23.73 51.50 41.12 43.53 41.64 40.33 53.19 28.12 35.27 64.25 43.80 39.57
ICon (75%) 15.17 23.79 19.48 55.11 29.80 46.70 39.94 42.89 39.99 37.24 52.52 28.73 35.12 64.25 42.98 39.03
@@
Table 6: The detailed results of different data scales on the evaluation benchmarks. The datasets AE, TQA, HS, LQA, and WG correspond to AlpacaEval, TruthfulQA, HellaSwag, LogicQA, and WinoGrande, respectively. All models are trained on LLaMA3.1-8B with 7,800 selected samples. The best results are highlighted in bold.
@@ Appendix C Detailed Evaluation Benchmark for Different Dataset
ICon (5%) 26.90 59.52 43.21 62.78 26.26 56.64 47.44 48.28 48.29 44.23 57.73 27.50 41.07 71.35 48.36 47.48 13.06
ICon (10%) 27.67 46.11 36.89 62.89 27.27 55.22 47.21 48.15 43.94 40.57 56.13 29.19 44.18 69.06 47.18 45.79 23.77
ICon (15%) 25.98 44.50 35.24 63.04 25.25 52.63 44.25 46.29 43.52 42.00 54.56 28.11 38.72 67.64 45.76 44.18 30.45
Table 7: The detailed performance of LLaMA3.1-8B trained on WizardLM data in terms of FLOPs and evaluation benchmarks for instruction following, knowledge, and reasoning. The datasets AE, TQA, HS, LQA, and WG correspond to AlpacaEval, TruthfulQA, HellaSwag, LogicQA, and WinoGrande, respectively. The best results of ICon across different training data scales are highlighted in bold.

### v2

@@ 7 Conclusion
In this paper, we propose RICo, a novel gradient-free data selection method that quantifies the refined contribution of individual samples for instruction tuning. RICo accurately identifies high-contribution samples by computing task- and global-level scores with reduced bias. We further introduce a lightweight selection paradigm, enabling efficient and scalable data curation. Experiments on three LLMs across 12 benchmarks and 5 pairwise evaluation sets show that models trained on RICo-selected data outperform full-data baselines, achieving better performance with fewer samples. RICo also outperforms other widely used selection methods. We additionally analyze optimal data scales, cross-dataset generalization, and the properties of selected samples.
@@
In estimating refined contribution, we leverage the implicit fine-tuning effect of in-context learning (ICL). Since ICL does not replicate the exact dynamics of full batch training with gradient updates, it may not fully capture all training interactions. However, it provides a lightweight and effective approximation of how individual samples influence model behavior, making it a practical tool for guiding data selection in instruction tuning. Furthermore, while our method demonstrates promising results for instruction tuning, we believe it has the potential for broader applicability. For instance, RICo could be adapted to other tasks and domains by modifying the assessment set and the calculation of the RICo score. Exploring these extensions is beyond the scope of this paper and remains a direction for future work.
@@ Appendix A Detailed Evaluation Benchmark for Different Methods
The detailed results for different methods on evaluation benchmarks are shown in Table 5. The methods include Low PPL, Top PPL, Alpagasus, Deita, Superfilter, and RICo. Overall, RICo outperforms existing methods across the evaluated benchmarks.
@@
Nuggets 13.78 23.48 18.63 53.95 30.81 51.69 41.80 44.56 43.09 44.18 54.92 27.50 40.28 67.32 46.22 41.07
RICo (Ours) 21.16 29.89 25.52 60.58 30.81 52.30 43.16 46.71 44.80 44.15 55.56 28.42 41.64 67.96 47.09 43.37
Table 5: The detailed results of different methods on the evaluation benchmarks. The datasets AE, TQA, HS, LQA, and WG correspond to AlpacaEval, TruthfulQA, HellaSwag, LogicQA, and WinoGrande, respectively. All models are trained on LLaMA3.1-8B with 15% selected samples.
@@ Appendix B Detailed Evaluation Benchmark for Different Data Scales
The detailed results for different data scales on evaluation benchmarks are shown in Table 6. As data scale increases, performance improves initially but later declines, indicating an optimal data size.
@@
RICo (1%) 8.31 28.62 18.46 58.47 31.31 55.77 43.11 47.17 49.91 44.51 61.19 23.35 39.37 71.03 48.23 42.91
RICo (5%) 16.21 28.96 22.58 56.91 27.27 53.70 40.51 44.60 47.18 44.23 58.44 24.42 37.75 69.46 46.91 42.09
RICo (10%) 16.08 27.99 22.03 57.86 26.76 53.57 42.63 45.21 45.48 42.81 56.27 26.57 42.01 66.93 46.68 42.08
RICo (15%) 21.16 29.89 25.52 60.58 30.81 52.30 43.16 46.71 44.80 44.15 55.56 28.42 41.64 67.96 47.09 43.37
RICo (20%) 11.95 27.24 19.60 55.07 29.29 53.02 41.71 44.77 43.36 42.08 55.06 26.88 37.36 64.88 44.94 40.66
RICo (25%) 15.63 23.35 19.49 50.84 28.28 51.12 41.26 42.88 42.06 42.13 54.52 28.88 39.77 67.32 45.78 40.43
RICo (30%) 17.72 25.87 21.79 58.57 24.74 52.41 40.66 44.10 41.89 42.32 54.35 26.57 37.63 66.54 44.88 40.77
RICo (50%) 15.00 22.91 18.95 57.75 23.73 51.50 41.12 43.53 41.64 40.33 53.19 28.12 35.27 64.25 43.80 39.57
RICo (75%) 15.17 23.79 19.48 55.11 29.80 46.70 39.94 42.89 39.99 37.24 52.52 28.73 35.12 64.25 42.98 39.03
@@
Table 6: The detailed results of different data scales on the evaluation benchmarks. The datasets AE, TQA, HS, LQA, and WG correspond to AlpacaEval, TruthfulQA, HellaSwag, LogicQA, and WinoGrande, respectively. All models are trained on LLaMA3.1-8B with 15% selected samples.
@@ Appendix C Detailed Evaluation Benchmark for Different Dataset
The detailed results of WizardLM dataset on evaluation benchmarks are shown in Table 7. As demonstrated by the experimental data, RICo exhibits strong generalization capabilities across diverse datasets.
@@
RICo (5%) 26.90 59.52 43.21 62.78 26.26 56.64 47.44 48.28 48.29 44.23 57.73 27.50 41.07 71.35 48.36 47.48 13.06
RICo (10%) 27.67 46.11 36.89 62.89 27.27 55.22 47.21 48.15 43.94 40.57 56.13 29.19 44.18 69.06 47.18 45.79 23.77
RICo (15%) 25.98 44.50 35.24 63.04 25.25 52.63 44.25 46.29 43.52 42.00 54.56 28.11 38.72 67.64 45.76 44.18 30.45
Table 7: The detailed performance of LLaMA3.1-8B trained on WizardLM data in terms of FLOPs and evaluation benchmarks for instruction following, knowledge, and reasoning. The datasets AE, TQA, HS, LQA, and WG correspond to AlpacaEval, TruthfulQA, HellaSwag, LogicQA, and WinoGrande, respectively.
Appendix D Additional References for Heuristic-based Methods
We list below all heuristic-based methods for automatic data selection in instruction tuning cited in Section 2:  Zhang et al. (2024b)Zhang et al. (2025),  Cao et al. (2023),  Chen et al. (2023),  Liu et al. (2023c),  Bukharin and Zhao (2023),  Minghao et al. (2024),  Du et al. (2023),  Kung et al. (2023),  Li et al. (2023a),  Li et al. (2024b),  Liu et al. (2024),  Wang et al. (2024),  Li et al. (2024a),  Ge et al. (2024),  He et al. (2024),  Han et al. (2025), and  Wettig et al. (2024).